OpenStudy (anonymous):

f(x)=x-1/x^2+1 g(x)=x+10/x+13 I need (f o g)(x) and (g o f)(x)

OpenStudy (anonymous):

(f o g) (x). put g(x) in f(x) as f(g(x)) Same with the other one. Just opposite. So it's f(g(x)) = (x+10/x+13) - 1 / (x+10/x+13) + 1 and g(f(x)) = (x-1/x^2+1) + 10 / (x-1/x^2+1) + 13

OpenStudy (anonymous):

thanks man